# Provably Overwhelming Transformer Models with Designed Inputs

### v1

@@ Appendix F Overwhelming for ${n_{ctx}}\to\infty$
We can note that $[X]=\vec{e}_{r}\cdot\text{E}$ E and $\mathrm{Var}[X]=\mathrm{Var}[\vec{e}_{r}\cdot\text{E}]=0$. So, we can easily compute blowup, $B$, and shift, $S$. Note, for attention head $\vec{e}_{n_{ctx}}\cdot f^{\texttt{AttnH}}_{(n_{fix},{n_{ctx}})\mid{s},q}$, the output equals
$\vec{p}(X)\cdot(XE\cdot\mathrm{diag}(B)+S)\cdot V$
for some probability vector $\vec{p}(X)$. Because $X=\vec{e}_{r}^{n_{ctx}}$, $(XE\cdot\mathrm{diag}(B)+S)\cdot V=\vec{a}^{n_{ctx}}$ for vector $\vec{a}=(\vec{e}_{r}\cdot E\cdot\mathrm{diag}(B)+S)V$. Thus,
$\vec{p}(X)\cdot(XE\cdot\mathrm{diag}(B)+S)\cdot V=\vec{a}.$
We therefore get that the output of $\vec{e}_{n_{ctx}}\cdot f^{\texttt{AttnH}}_{(n_{fix},{n_{ctx}})\mid{s},q}$ is fixed to $\vec{e}_{r}(\text{E}\cdot\mathrm{diag}(B)+S)\cdot V$ for all context windows ${n_{ctx}}$. Finally, we can simply compute the value of $\vec{e}_{n_{ctx}}\cdot f^{\text{MLP}}_{(n_{fix},{n_{ctx}})\mid{s},q}$ because we know $B$ and $S$ and so just need to compute $\text{MLP}(\vec{e}_{r}(E\cdot\mathrm{diag}(B)+S))$. ∎
Bounds on Expectation and Variance
Lemma F.4 (Bounds on Expectation and Variance).
Let $X^{\prime}\in\mathbb{O}_{rep}$ and $X=\mathopen{}\mathclose{{}\left(\bigtimes_{i=1}^{n_{fix}}\vec{e}_{r}}\right)\times X^{\prime}\times\vec{e}_{q}$. I.e. $X$ is an expansion of the input with the first tokens being the repetition token and the last being the query token. Then, we can bound the column wise expectation for column $j$, $\mu_{j}$, and variance, $\mathrm{Var}_{j}$ of $X$ by
$\frac{1}{{n_{ctx}}}\mathopen{}\mathclose{{}\left(n_{fix}\cdot\vec{e}_{r}\text{E}\vec{e}_{j}^{T}+n_{free}\min_{i}(\vec{e}_{i}\text{E}\vec{e}_{j}^{T})+\vec{e}_{q}\text{E}\vec{e}_{j}^{T}}\right)\leq\mu_{j}\leq$
$\frac{1}{{n_{ctx}}}\mathopen{}\mathclose{{}\left(n_{fix}\cdot\vec{e}_{r}\text{E}\vec{e}_{j}^{T}+n_{free}\max_{i}(\vec{e}_{i}\text{E}\vec{e}_{j}^{T})+\vec{e}_{q}\text{E}\vec{e}_{j}^{T}}\right).$
Let $\mu_{j}^{\min}$ and $\mu_{j}^{\max}$ denote the lower and upper bounds. Then, we can bound the variance of the $j$-th column of $\vec{x}\text{E}$ by
$\mathrm{Var}_{j}\leq$ $\frac{1}{{n_{ctx}}}\cdot\max_{\mu^{\prime}\in[\mu_{j}^{\min},\mu_{j}^{\max}]}\bigg{[}(n_{fix}+1)\cdot\mathopen{}\mathclose{{}\left(\vec{e}_{r}\text{E}\vec{e}_{j}^{T}-\mu^{\prime}}\right)^{2}+n_{free}\cdot\max_{i\in{d_{vocab}}}\mathopen{}\mathclose{{}\left(\vec{e}_{i}\text{E}\vec{e}_{j}^{T}-\mu^{\prime}}\right)^{2}\bigg{]}.$
We define $\mathrm{Var}_{j}^{\min},\mathrm{Var}_{j}^{\max}$ to be the lower and upper bounds respectively.
Proof.
The bounds on $\mu_{j}$ follow as we minimize and maximize the contribution of each free row to the expectation of the column of $j$. The proof of for variance bounds follows similarly. For the lower bound, we note that the variance contribution from each row is at least 00. For the upper bound, we simply maximize the contribution of each row to the variance. ∎
Recall the definition of blowup and shift sets from definition 4.6.
Lemma F.5 (Blowup and Shift Set Bounds).
Recall that $\gamma$ and $\beta$ are the learned constants for layer normalization. Let $\vec{x}\in\mathbb{O}_{des}$. Then, we can bound the blowup and shift sets for the $j$-th column of $\vec{x}\text{E}$. For all $S_{j}\in\mathcal{S}_{j}$ and $B_{j}\in\mathcal{B}_{j}$, we have
$\frac{\gamma}{\sqrt{\mathrm{Var}_{j}^{\max}+\epsilon}}\leq B_{j}\leq\frac{\gamma}{\sqrt{\epsilon}}$
Let $B_{j}^{\min}$ and $B_{j}^{\max}$ denote the lower and upper bounds respectively. Then,
$B_{j}^{\min}\mu_{j}^{\min}+\beta\leq S_{j}\leq B_{j}^{\max}\mu_{j}^{\max}+\beta.$
Now as a corollary of lemma F.4 and lemma F.5, we can show that the minimum and maximums of the blowup and shift sets can be made arbitrarily close together.
Corollary F.6 (Convergence of Blowup and Shift Sets).
If $\frac{n_{free}}{{n_{ctx}}}\in o(1)$, then for every $\delta_{1},\delta_{2}\in(0,1)$, there exists some setting of ${n_{ctx}}^{\prime}$ such that for all ${n_{ctx}}>{n_{ctx}}^{\prime}$,
$B_{j}^{\max}-B_{j}^{\min}\leq\delta_{1}\quad\text{and}\quad S_{j}^{\max}-S_{j}^{\min}\leq\delta_{2}.$
Proof.
The corollary follows from the fact that $\mu_{j}^{\min},\mu_{j}^{\max}$ converge to $\vec{e}_{r}E\vec{e}_{j}^{T}$ as ${n_{ctx}}\rightarrow\infty$. So then, $(\vec{e}_{r}E\vec{e}_{j}^{T}-\mu^{\prime})^{2}$ for all $\mu^{\prime}\in[\mu_{j}^{\min},\mu_{j}^{\max}]$ converges to 00 and thus the upper bound on $\mathrm{Var}_{j}$ converges to 00 as long as $\frac{n_{free}}{{n_{ctx}}}$ converges to 00. ∎
Therefore the blowup and shift sets converge, let $B_{j}^{\prime}$ and $S_{j}^{\prime}$ denote the converged values for $B_{j}$ and $S_{j}$ respectively.
Now, we need to show a bound on the difference between the maximum logit for the free tokens and the minimum logit for the repeated tokens in the attention head.
@@
Lemma F.8 (Bound on Difference).
We can bound the difference between the maximum logit for the free tokens and the minimum logit for $n_{fix}$ repeated tokens
$L_{free}^{\max}-L_{fix}^{\min}$ $\leq\theta$ (5)
and
$L_{free}^{\min}-L_{fix}^{\max}$ $\leq\theta$ (6)
where $\theta$ is a small constant.
Proof.
We will prove the first inequality as the second follows analogously. Assume that $i\in[{d_{vocab}}]$ is the token which maximizes the difference. We will relly on the convergence of the blowup and shift sets to show that the difference between the maximum logit for the free tokens and the minimum logit for the repeated tokens in the attention head is bounded.
Connote $\mathcal{B}-B^{\prime}$ as the set of blowup values shifted by the converged value and $\mathcal{S}-S^{\prime}$ as the set of shift values shifted by the converged value. Then $\mathcal{B}-B^{\prime}$ and $\mathcal{S}-S^{\prime}$ are bounded by $\delta_{1},\delta_{2}$ respectively, we can write the following bound on the difference.
$L_{free}^{\max}-L_{fix}^{\min}$ $\leq\max_{\epsilon\in\mathcal{B}-B^{\prime},\;\epsilon_{S}\in\mathcal{S}-S^{\prime}}(\vec{e}_{i}-\vec{e}_{r})E\cdot\mathrm{diag}(B^{\prime}+\epsilon)$
$\cdot QK^{T}(\mathrm{diag}(B^{\prime}+\epsilon)\cdot E^{T}\vec{e}_{q}^{T}+S^{T}+\epsilon_{S})$
$=\max_{\epsilon\in\mathcal{B}-B^{\prime},\epsilon_{S}\in\mathcal{S}-S^{\prime}}$
$\sum_{d_{1}\in\{\epsilon,B^{\prime}\}}\sum_{d_{2}\in\{\epsilon,B^{\prime}\}}\sum_{d_{3}\in\{\epsilon,S^{\prime}\}}(\vec{e}_{i}-\vec{e}_{r})E\cdot\mathrm{diag}(d_{1})\cdot QK^{T}(\mathrm{diag}(d_{2})\cdot E^{T}\vec{e}_{q}^{T}+d_{3})$
$\leq\max_{\epsilon\in\mathcal{B}-B^{\prime},\epsilon_{S}\in\mathcal{S}-S^{\prime}}(\vec{e}_{i}-\vec{e}_{r})E\cdot\mathrm{diag}(B^{\prime})\cdot QK^{T}(\mathrm{diag}(B^{\prime})\cdot E^{T}\vec{e}_{q}^{T}+S^{T})+{\mathsf{poly}}(\epsilon,\epsilon_{S})$
$\leq(\vec{e}_{i}-\vec{e}_{r})E\cdot\mathrm{diag}(B^{\prime})\cdot QK^{T}(\mathrm{diag}(B^{\prime})\cdot E^{T}\vec{e}_{q}^{T}+S^{T})+{\mathsf{poly}}(\delta_{1},\delta_{2},B^{\prime},S^{\prime})$ (By the bounds on the blowup and shift set)
$\leq(\vec{e}_{i}-\vec{e}_{r})E\cdot\mathrm{diag}(B^{\prime})\cdot QK^{T}(\mathrm{diag}(B^{\prime})\cdot E^{T}\vec{e}_{q}^{T}+S^{T})+{\mathsf{poly}}(\delta_{1},\delta_{2})$
where the last inequality follows from considering $B^{\prime}$ and $S^{\prime}$ as a constant. Then, because $\delta_{1},\delta_{2}$ decrease as a function of ${n_{ctx}}$ and corollary F.7, we can bound
$(\vec{e}_{i}-\vec{e}_{r})E\cdot\mathrm{diag}(B^{\prime})\cdot QK^{T}(\mathrm{diag}(B^{\prime})\cdot E^{T}\vec{e}_{q}^{T}+S^{T})+{\mathsf{poly}}(\delta_{1},\delta_{2})\leq\theta$
for a large enough ${n_{ctx}}$. ∎
@@
$\mathcal{W}(\vec{e}_{{n_{ctx}}}\cdot f^{\texttt{AttnH}}_{n_{fix}\mid r,q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}\leq({\beta}_{fix}^{\max}-{\beta}_{fix}^{\min})\cdot\max_{B,S}\norm{\vec{e}_{r}\text{E}B+S}+2\cdot{\beta}_{free}^{\max}\cdot\max_{B,S}\norm{(\vec{e}_{r}\text{E}B+S)\cdot V}_{\infty}$
@@
$\mathcal{W}(f^{\texttt{AttnH}}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des}\times\mathcal{B}\mathcal{S})_{\infty}\leq o(1)$
@@
Bound on $f^{enc}$
Finally, as per section 4, we need to bound the worst-case deviation of the $f^{enc}$ function. Given that $f^{enc}$ is a linear operation as a function of blowup and shift sets, we can use the convergence of the blowup and shift sets to show that the worst-case deviation of $f^{enc}$ is bounded.
Lemma F.10 (Bound on $f^{enc}$).
For large enough ${n_{ctx}}$, we have
$\mathcal{W}(\vec{e}_{n_{ctx}}\cdot f^{enc}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{rep})_{\infty}\leq{\mathsf{poly}}(\delta_{1},\delta_{2}).$
, where $\delta_{1},\delta_{2}$ are the bounds on the blowup and shift differences.
@@
Note that we reduced $\mathcal{W}(\vec{e}_{n_{ctx}}\cdot\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{rep})_{\infty}$ to be upper-bounded by $O(1)\cdot\mathcal{W}(\vec{e}_{n_{ctx}}\cdot f^{enc}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{rep})+o(1)$ through lemma 4.14 and lemma F.9. Then, by lemma F.10, we have that theorem F.2 holds as $\mathcal{W}$ goes to 00 as ${n_{ctx}}\rightarrow\infty$. So, as long as $\mathrm{PTP}(\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q},X)$ is positive for some $X\in\mathbb{O}_{rep}$, then we converge to “overwhelming” by theorem 4.4. Note that $\vec{e}_{r}^{n_{ctx}}\in\mathbb{O}_{rep}$ and by lemma F.3, we can compute a sample of peak-to-peak deviation for all ${n_{ctx}}$. ∎

### v2

@@ Appendix F Overwhelming for ${n_{ctx}}\to\infty$
So, we can easily compute blowup, $B$, and shift, $S$. Note, for attention head $\vec{e}_{n_{ctx}}\cdot f^{\texttt{AttnH}}_{(n_{fix},{n_{ctx}})\mid{s},q}$, the output equals
$\vec{p}(X)\cdot X\text{E}_{\texttt{LN}}\cdot V$
for some probability vector $\vec{p}(X)$. Because $X=\vec{e}_{r}^{n_{ctx}}$, $X\text{E}_{\texttt{LN}}\cdot V=\vec{a}^{n_{ctx}}$ for vector $\vec{a}=(\vec{e}_{r}\cdot\text{E}_{\texttt{LN}})V$. Thus,
$\vec{p}(X)\cdot(X\text{E}_{\texttt{LN}})\cdot V=\vec{a}.$
We therefore get that the output of $\vec{e}_{n_{ctx}}\cdot f^{\texttt{AttnH}}_{(n_{fix},{n_{ctx}})\mid{s},q}$ is fixed to $\vec{e}_{r}\text{E}_{\texttt{LN}}\cdot V$ for all context windows ${n_{ctx}}$. Finally, we can simply compute the value of $\vec{e}_{n_{ctx}}\cdot f^{\text{MLP}}_{(n_{fix},{n_{ctx}})\mid{s},q}$ because we know $B$ and $S$ and so just need to compute $\text{MLP}(\vec{e}_{r}\text{E}_{\texttt{LN}})$. ∎
@@
$\mathcal{W}(\vec{e}_{{n_{ctx}}}\cdot f^{\texttt{AttnH}}_{n_{fix}\mid r,q};\mathbb{O}_{des})_{\infty}\leq({\beta}_{fix}^{\max}-{\beta}_{fix}^{\min})\cdot\norm{\vec{e}_{r}\text{E}_{\texttt{LN}}}+2\cdot{\beta}_{free}^{\max}\cdot$
$norm{(\vec{e}_{r}\text{E}_{\texttt{LN}})\cdot V}_{\infty}$
@@
$\mathcal{W}(f^{\texttt{AttnH}}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{des})_{\infty}\leq o(1)$
@@
Note that we reduced $\mathcal{W}(\vec{e}_{n_{ctx}}\cdot\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{rep})_{\infty}$ to be upper-bounded by $O(1)\cdot\mathcal{W}(\vec{e}_{n_{ctx}}\cdot f^{enc}_{(n_{fix},{n_{ctx}})\mid{s},q};\mathbb{O}_{rep})+o(1)$ through lemma 4.12 and lemma F.5. Then, by theorem 4.5, we have that theorem F.2 holds as $\mathcal{W}$ goes to 00 as ${n_{ctx}}\rightarrow\infty$. So, as long as $\mathrm{PTP}(\mathcal{M}_{(n_{fix},{n_{ctx}})\mid{s},q},X)$ is positive for some $X\in\mathbb{O}_{rep}$, then we converge to “overwhelming” by theorem 4.4. Note that $\vec{e}_{r}^{n_{ctx}}\in\mathbb{O}_{rep}$ and by lemma F.3, we can compute a sample of peak-to-peak deviation for all ${n_{ctx}}$. ∎